OpenStudy (anonymous):

4x+2y=8 4x+3y=16 x = ? y = ?

OpenStudy (ranga):

Subtract the first equation from the second and you will get the y value. Put y in the first equation and solve for x.

OpenStudy (anonymous):

so what x = 8?

OpenStudy (ranga):

y = 8

OpenStudy (anonymous):

so y = 8 than x = ? x = 8

OpenStudy (ranga):

Put y = 8 in the first equation: 4x+2y=8 4x + 2(8) = 8 4x + 16 = 8 4x = -8 x = -2

OpenStudy (anonymous):

thanks can u help me with the next

OpenStudy (ranga):

I can assist you but try to learn the method and then you don't have to ask anyone and you can do it much faster yourself.